our galaxy is about 15 light years across and most energetic particles transversing it have energies of about 10^20 ev find the relativistic energy of the partical assuming it in a proton
1)16J
2)15J
3)10J
4)0J
1
Expert's answer
2019-04-08T09:37:25-0400

So, we have energy

"E=10^{20} eV"

for one eV

"1 eV=1.6\cdot 10^{-19} J"

In joules this energy

"E=10^{20} \cdot 1.6 \cdot 10^{-19} = 16 J"

Answer 1) 16 J